The indentfirst package^{∗}

David Carlisle carlisle@cs.man.ac.uk

1995/11/23

Abstract

Make the first line of all sections etc., be indented by the usual paragraph indentation. This should work with all the standard document classes.

\if@afterindent L^{A}TEX uses the switch \if@afterindent to decide whether to indent after a section heading. We just need to make sure that this is always true.

1 ⟨∗package⟩

2 \let\@afterindentfalse\@afterindenttrue
3 \@afterindenttrue

4 ⟨/package⟩

∗This file has version number v1.03, last revised 1995/11/23.